Solve.
9x^4-18x^2+5=0



Answer by ewatrrr(24785)

 
Hi,
9x^4-18x^2+5=0
factoring
(3x^2 - 1)(3x^2 - 5) = 0
(3x^2 - 1)= 0 x = sqrt(1/3)
(3x^2 - 5) = 0 x = sqrt(5/3)